OpenStudy (anonymous):

Can anyone explain to me how to write the equation of a line through C that is parallel to line AB. The slope of line AB is 3/4. The slope of line AC is -1/7. On a graph, A is (-2,0), B is (2,3), and C is (5,-1).

OpenStudy (precal):

take the slope of AB and the point of C m=3/4 x=5 y=-1 and use y=mx+b solve for b

OpenStudy (precal):

by the slope you already know that your line is y=(3/4)x+b it is the b value that you are missing